5g of NaOH and NaCL ( as impurity) were dissolved in 1dm3 of water. If 25cm3 of this solution were neutralized by 21cm3 of 0.1moldm3 HCL. Calculate the mass concentration of NaOH in gdm3.


1
Expert's answer
2022-05-11T13:53:03-0400

NaOH + HCl ==> NaCl + H2O

Converting volumes from cm3 to dm3:

volume of HCl = 21.0 cm3 = 21.0 ÷ 1000 = 0.021 dm3

volume of NaOH = 25.0 cm3 = 25.0 ÷ 1000 = 0.025 dm3


Amount of HCl in mol = concentration in mol/dm3 × volume in dm3

= 0.1 mol/dm3 × 0.021 dm3

= 0.0021 mol

The equation shows that 1 mol of HCl reacts with 1 mol of NaOH. So the amount of NaOH in 25.0 cm3 is 0.0021 mol.

[NaOH] = 0.0021 mol / 0.025 dm3 = 0.084 mol/dm3